solve the equation
|x^2+2x-36|=12

Thanks

Answer by stanbon(57214)

x^2+2x-36=12 or x^2+2x-36=-2
x^2+2x-48=0 or x^2+2x-34=0
(x+8)(x-6)=0 or x=[-2+-sqrt(4-4(-24)]/2
x=-8 or x=6 or x=[-1+-4sqrt6]
x=-8 or x=6 or x=-1+4sqrt6 or x=-1-4sqrt6
Cheers,
Stan H.